Circular No. 8816

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2004 PB_{108} AND (60621) 2000 FE_{8}

K. S. Noll, Space Telescope Science Institute (STScI); W. M. Grundy, Lowell Observatory; S. D. Kern, STScI; H. F. Levison, Southwest Research Institute; and D. C. Stephens, Brigham Young University, report the detection of binary companions to two additional transneptunian objects: 2004 PB_{108} (cf. MPECs 2004-R17; MPO 86538) and (60621) 2000 FE_{8} (cf. MPECs 2000-F46; MPO 110654). Both were found on images obtained with the High Resolution Camera of the Advanced Camera for Surveys on the Hubble Space Telescope, using the clear filters with one 300-s exposure at each of four dithered positions on the detector. The observations of 2004 PB_{108} were made during 2006 Aug. 4.300–4.377 UT; its two components were separated by an angular distance of 0′′.172 ± 0′′.003 and differ in brightness by 1.2 mag, with the fainter component lying at a position angle of 298^{o} ± 2^{o} from the primary (the projected separation of the objects in the sky plane is 5320 ± 90 km). The observations of (60621) were made during 2007 Jan. 15.066–15.091; its two components were separated by an angular distance of 0′′.044 ± 0′′.003 and differ in brightness by 0.6 mag, the fainter component lying at a position angle of 187^{o} ± 11^{o} from the primary (the projected separation of the objects in the sky plane is 1180 ± 80 km).

COMET C/2006 P1 (McNAUGHT) N. Dello Russo, R. J. Vervack, Jr., H. A. Weaver, and C. M. Lisse,

Applied Physics Laboratory, Johns Hopkins University, report the detection of parent volatiles in comet C/2006 P1 on Jan. 27 (r = 0.55 AU, Δ = 1.05 AU, geocentric velocity = +38.2 km/s) using the NASA Infrared Telescope Facility (+ CSHELL). The following species were detected: H_{2}O, CO, CH_{4}, HCN, C_{2}H_{2}, C_{2}H_{6}, NH_{3}, H_{2}CO, CH_{3}OH, NH_{2}, and OH. For an aperture of size 1′′ × 3′′ centered on the nucleus, they obtained the following line fluxes (× 10^{−18} W m^{−2}) for representative emission lines: H_{2}O (ν_{3}-ν_{2} 2_{20}-2_{21}), 60; CO (1-0 R0), 20; CH_{4} (ν_{3} R0), 9.8; HCN (ν_{3} R6), 19; C_{2}H_{2} (ν_{3} R3), 15; C_{2}H_{6} (ν_{7} ^{R}Q_{0}), 38; NH_{3} [ν_{1} aqQ(3,3)], 10; H_{2}CO (ν_{1}, several blended lines at 2781 cm^{−1}), 11. The continuum flux densities [× 10^{−15} W m^{−2} (cm^{−1})^{−1}] within the above aperture were 1.2, 1.7, 2.1, and 7 at 3.02, 3.30, 3.59, and 4.69 microns, respectively. Preliminary results are given assuming a rotational temperature of 150 K for all species. The H_{2}O production rate was 1.7 × 10^{30} molecules/s, and the abundances relative to H_{2}O were as follows: CO, 2 percent; CH_{4}, 0.45 percent; C_{2}H_{6}, 0.45 percent; C_{2}H_{2}, 0.5 percent; H_{2}CO, 0.5 percent; NH_{3}, 2 percent; HCN, 0.25 percent.

2007 March 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green